OpenStudy (anonymous):

Simplify the following expression. 2x - 10y + 3x2 + 7y - 10x.

OpenStudy (anonymous):

\(2x-10x=-8x\) since \(2-10=-8\) what is \(7y-10y\) ?

OpenStudy (anonymous):

3y @satellite73

OpenStudy (anonymous):

close

OpenStudy (anonymous):

\(7-10=-3\) though, not \(3\)

OpenStudy (anonymous):

@satellite73 oh..

OpenStudy (anonymous):

so you have \[2x-10x=-8x\\7y-10y=-3y\] and \(3x^2\) just hanging out by itself final answer will be \[3x^2-8x-3y\] in any order you choose

OpenStudy (anonymous):

@satellite73 thank you!